# Efficient Inference for Large Reasoning Models: A Survey

### v1

@@ 6.3 Agent Router
There are two existing routing strategies: one based on router models and the other on metrics like confidence. Routellm (Ong et al., 2025) introduces several efficient router models that dynamically select either a stronger or a weaker LLM during inference to balance cost and response quality. Self-REF (Chuang et al., 2025b) routes LLMs by training them to reliably express their confidence in the correctness of their answers. Additionally, Chuang et al. (2025a) explores an uncertainty-based small language model (SLM) routing approach, where high-stakes queries are offloaded to more robust LLMs whenever the SLM produces low-confidence responses. These methods offer exciting possibilities for enhancing inference efficiency in LRMs, particularly in reducing computation time and resource use while maintaining high performance across a range of tasks.

### v2

@@ 6.2 Model Merge
Merging models with disparate architectures or significantly different layer counts presents a substantial technical hurdle. Simple averaging is often not viable. This may require more sophisticated techniques like parameter subspace alignment, or focusing the merge on specific, architecturally compatible layers (e.g., only attention or feed-forward layers if they share dimensions), which might limit the extent of synergistic merging. Compared to other ensemble methods (like distillation or mixture-of-experts where models are used more distinctly), model merging aims for a deeper integration at the parameter level. However, its practical advantage over simpler ensembling or parameter-efficient fine-tuning (PEFT) methods applied to a single base model needs to be clearly demonstrated, especially concerning the development effort and the consistency of performance gains.
@@ 6.3 Agent Router
There are two existing routing strategies: one based on router models and the other on metrics like confidence. Routellm (Ong et al., 2025) introduces several efficient router models that dynamically select either a stronger or a weaker LLM during inference to balance cost and response quality. Self-REF (Chuang et al., 2025b) routes LLMs by training them to reliably express their confidence in the correctness of their answers. Additionally, Chuang et al. (2025a) explores an uncertainty-based small language model (SLM) routing approach, where high-stakes queries are offloaded to more robust LLMs whenever the SLM produces low-confidence responses. These methods offer exciting possibilities for enhancing inference efficiency in LRMs, particularly in reducing computation time and resource use while maintaining high performance across a range of tasks. Comparatively, agent routing needs to demonstrate clear advantages over simpler strategies like multi-pass inference with a single scalable model (e.g., an initial quick generation followed by a more deliberative self-correction or refinement step if the initial output is low-confidence). While routing to specialized agents appears compelling for diverse tasks, the practicalities of developing, maintaining, and orchestrating a suite of specialized models, along with a robust router, are substantial. This includes ensuring consistent API interactions, managing different model versions, and updating the router’s logic as new agents are added or existing ones are modified. The complexity of the overall system increases, which can impact deployment and operational costs. The true benefit emerges when the specialization of agents and the accuracy of routing lead to a system that is significantly more efficient or capable than a single, larger general-purpose model or simpler adaptive computation schemes.